cos2+sin2=?

cos2+sin2=?

Grade:11

2 Answers

Vijay Mukati
8 years ago
Dear Student,

Please write the question clearly.

Thanks.
arun
8 years ago
see i think you want to simlify \sin 2+\cos2
now, \sin 2+\cos2 = \sin2 + \cos(90-88)
\Rightarrow \sin 2+\cos2 = \sin2 + \sin88
\Rightarrow \sin 2+\cos2 = 2\sin(\frac{88+2}{2})\cos(\frac{88-2}{2})
\Rightarrow \sin 2+\cos2 = 2\sin(45)\cos(43)
\Rightarrow \sin 2+\cos2 = \sqrt{2}\cos(43) 
here is your answer that sin2 + cos2 = (square root of 2) cos43
you ask me if your question was different, there is any mistake in my answer or there is any doubt related to this question